STAT 230 Chapter 2, Spring 2020
American University of Beirut · Introduction to Probability and Random Variables · 82 pages
Showing the first 10 of 82 pages. Download the PDF to see all of it.
More STAT 230 files
- Chapter 2 · Spring 2020ChapterView
- Syllabus · Spring 2020LabView
- Useful File Solution Manual · Spring 2019File · with solutionView
- Lecture 3- Chapters 2.1, 2.2 & 2.4 · Spring 2019LabView
- Lecture 1- Chapter 1 · Spring 2019LabView
- Lecture 2- Chapters 2.1 & 2.2 · Spring 2019LabView
- Lecture 4- Examples About Chapters 2.4 · Spring 2019LabView
- Lecture 4- Chapter 2.4 · Spring 2019LabView
Show the text of this chapter
Chapter 2- 2020 Summer
A Set-theory Recap
Let A, B be two subsets of a set Ω
Union
A ∪ B = {ω ∈ Ω : ω ∈ A or ω ∈ B}
Intersection
A ∩ B ≡ AB = {ω ∈ Ω : ω ∈ A and ω ∈ B}
Complement
¯A ≡ Ac = {ω ∈ Ω : ω /∈ A} = Ω \ A
Difference
B \ A = B ∩ ¯A = {ω ∈ Ω : ω ∈ B and ω /∈ A}
Notation
When A is a countable set, we denote with |A| the number of
points/elements in A (the size or cardinality of A)
A Set-theory Recap
Properties
Commutativity
A ∪ B = B ∪ A,
A ∩ B = B ∩ A
Associativity
(A ∪ B) ∪ C = A ∪ (B ∪ C),
(A ∩ B) ∩ C = A ∩ (B ∩ C)
Distributivity
A∩(B∪C) = (A∩B)∪(A∩C),
A∪(B∩C) = (A∪B)∩(A∪C)
Idempotency
A ∪ A = A,
A ∩ A = A
De Morgan’s laws
A ∪ B = ¯A ∩ ¯B,
A ∩ B = ¯A ∪ ¯B
n�
i=1
Ai =
n�
i=1
¯Ai,
n�
i=1
Ai =
n�
i=1
¯Ai
A Set-theory Recap. Poincar´e Identities
Inclusion-Exclusion Principle (Poincar´e Identity)
Let Ai ⊂ Ω, i = 1, . . . , n
|
n�
i=1
Ai|
=
|A1 ∪ A2 ∪ . . . ∪ An| =
n
�
i=1
|Ai| −
�
1≤i<j≤n
|Ai ∩ Aj|
+
�
1≤i<j<k≤n
|Ai ∩ Aj ∩ Ak|
− . . . + (−1)n−1|A1 ∩ A2 ∩ . . . An|
For example, for n = 3
|A∪B ∪C| = |A|+|B|+|C|−|A∩B|−|A∩C|−|B ∩C|+|A∩B ∩C|
A Set-theory Recap. Poincar´e Identities
Using the De Morgan’s law, it then follows
|
n�
i=1
¯Ai|
=
|Ω \
n�
i=1
Ai|
=
|Ω| − |
n�
i=1
Ai|
since Ω ⊇
n�
i=1
Ai
=
|Ω| −
n
�
i=1
|Ai| +
�
1≤i<j≤n
|Ai ∩ Aj|
−
�
1≤i<j<k≤n
|Ai ∩ Aj ∩ Ak|
+ . . . + (−1)n−1|A1 ∩ A2 ∩ . . . An|
An Example
In a group of 21 (Arabic-speaking) Lebanese, 16 speak French, 13
English, 4 Armenian, 9 English and French, 2 French and
Armenian, 3 English and Armenian and 1 English, French and
Armenian. How many speak only Arabic?
An Example
In a group of 21 (Arabic-speaking) Lebanese, 16 speak French, 13
English, 4 Armenian, 9 English and French, 2 French and
Armenian, 3 English and Armenian and 1 English, French and
Armenian. How many speak only Arabic?
Let A be the set of Armenophones, E the set of Anglophones, F
the set of Francophones
The set of those who speak only Arabic is
An Example
In a group of 21 (Arabic-speaking) Lebanese, 16 speak French, 13
English, 4 Armenian, 9 English and French, 2 French and
Armenian, 3 English and Armenian and 1 English, French and
Armenian. How many speak only Arabic?
Let A be the set of Armenophones, E the set of Anglophones, F
the set of Francophones
The set of those who speak only Arabic is
¯A ∩ ¯E ∩ ¯F
=
A ∪ E ∪ F = Ω \ (A ∪ E ∪ F)
| ¯A ∩ ¯E ∩ ¯F|
=
|Ω| − |A ∪ E ∪ F|
=
|Ω| − (|A| + |E| + |F|) + |A ∩ E| + |A ∩ F| + |E ∩ F|
−|A ∩ E ∩ F|
=
21 − (4 + 13 + 16) + 3 + 2 + 9 − 1 = 1
Sample Space and Events
The sample space Ω (or S) associated with an experiment is the
set of all possible outcomes of such an experiment
Ω = {ω1, . . . , ωn}
A subset E of Ω, E ⊂ Ω, is called an event
Informally, an event is a statement on the outcomes of a random
experiment
We also assume that
Ω is an event (the certain or universal event)
if E is an event, so is its complement ¯E
Hence, the empty set ∅ is an event (the impossible, or
vacuous, event)
the union of events is an event
Sample Space and Events
When E is just one outcome we say it is a simple event or a state
(e.g. E = {ω3})
When E is more than one, we say it is a composite event or,
simply, event (e.g. E = {ω2, ω21})
Sample Space
Describe the sample spaces for the following experiments
▶ roll a regular die
Sample Space
Describe the sample spaces for the following experiments
▶ roll a regular die
{1, 2, 3, 4, 5, 6}
▶ toss n coins (n = 4)
Probability in a Simulation
Roll a die, what is the probability of A =”3 appears”?
The pseudo-code for the simulation to get the probability P(A)
Problem
How many ways can we roll three dice? In how many ways can
three dice appear when they are rolled? How many possible
numbers we get by rolling 3 dice (the order counts)?
Some problems
▶ How many ways are there for 2 persons to sit in 7 chairs that
are in a row?
7 · 6 = (7)2
Sampling (the chairs) with ordering (it matters in a theater,
say, where you sit) and without replacement (two people may
not sit on the same chair)
▶ How many ways are there to make a 3-letter word (a string of
3 letters) using a 26-letter alphabet?
263
▶ How many ways are there to form a 3-letter word if the letters
must be different?
Sample Space
Describe the sample spaces for the following experiments
▶ roll a regular die
{1, 2, 3, 4, 5, 6}
▶ toss n coins (n = 4)
{HHHH, HHHT, HHTH, HHTT, HTHH, HTTH, HTHT, HTTT,
TTTT, TTTH, TTHT, TTHH, THTT, THHT, THTH, THHH}
▶ cast a regular die and, if 6 comes up, toss a coin
Sample Space
Describe the sample spaces for the following experiments
▶ roll a regular die
{1, 2, 3, 4, 5, 6}
▶ toss n coins (n = 4)
{HHHH, HHHT, HHTH, HHTT, HTHH, HTTH, HTHT, HTTT,
TTTT, TTTH, TTHT, TTHH, THTT, THHT, THTH, THHH}
▶ cast a regular die and, if 6 comes up, toss a coin
6H, 6T, 1, 2, 3, 4, 5
▶ measure the time for the emission of radioactive particle from
some atom
Sample Space
Describe the sample spaces for the following experiments
▶ roll a regular die
{1, 2, 3, 4, 5, 6}
▶ toss n coins (n = 4)
{HHHH, HHHT, HHTH, HHTT, HTHH, HTTH, HTHT, HTTT,
TTTT, TTTH, TTHT, TTHH, THTT, THHT, THTH, THHH}
▶ cast a regular die and, if 6 comes up, toss a coin
6H, 6T, 1, 2, 3, 4, 5
▶ measure the time for the emission of radioactive particle from
some atom
(0, ∞)
(this is a non-discrete case)
Events
Describe the events
1) when rolling a die, an even number comes up
2) when tossing n coins, the first n − 1 outcomes are tails
3) measuring the time for the emission of radioactive particle
from some atom, the emission occurs after 3 minutes
Events
Describe the events
1) when rolling a die, an even number comes up
2) when tossing n coins, the first n − 1 outcomes are tails
3) measuring the time for the emission of radioactive particle
from some atom, the emission occurs after 3 minutes
1) E = {2, 4, 6}
2)
E = {
n−1
�
��
�
TTTTTT T,
n−1
�
��
�
TTTTTT H}
3) E = (3, ∞) (using minutes as unit of time)
Sample Space and Events
Events are subsets of a set, the sample space Ω
Logical statements on events ⇐⇒ operation with sets
Both event A and event B occur (conjunction) ⇐⇒ A ∩ B
at least one of A and B occurs (disjunction)/ either A or B or
both occur ⇐⇒ A ∪ B
the event A does not occur (negation) ⇐⇒
¯A
the event A occurs but the event B does not ⇐⇒ A \ B
Some Terminology
If A ∩ B = ∅, A and B are said to be disjoint or incompatible
or mutually exclusive events
(that A and B both occur is impossible: the occurrence of one
prevents the occurrence of the other)
If A ⊆ B, A is said to imply B (B occurs if A occurs)
In fact ⊆ is the set-theoretic equivalent of ⇒
Probability
The first two elements of a probabilistic model are the sample
space and the notion of events.
The third element is the assignment of a probability to the events
and outcomes of a random experiment
We need to formalize statements such as
the probability that when we roll a die an even number comes up is
1/2
Probability as Measure of Relative Frequencies
One interpretation views probability as a relative frequency (which
can be justified a posteriori by the result known as the law of large
numbers)
Carry out repeatedly and independently the same experiment a
large number of times N (roll the same die in the same conditions
N times)
record the number of times SN(E) the event E occurs (”an even
number comes up”)
assign to the event the probability P(E) = SN(E)/N, N large, (the
empirical limiting relative frequency in the N repetitions)
Probability as Measure of Relative Frequencies
The frequency definition of probability is based on the assumption
that identical and independent experiments can be carried out
(which is not always the case)
A priori there is no guarantee that the relative frequency should
converge to a limit and if that is the case, it is not clear how large
N should be for the approximation to be reliable
However even if there may be difficulties involved in defining
probability in a mathematical form using repetitive events, this
notion of probability is the basis of simulations, so you should keep
it in mind
Probability in a Simulation
Roll a die, what is the probability of A =”3 appears”?
The pseudo-code for the simulation to get the probability P(A)
counter_a=0
for i from 1 to N
roll a die
if(die
shows 3)
counter_a = counter_a+1
end if
end for
return counter_a/N
Probability in a Simulation. Implementation
▶ Python
import numpy as np
np.random.seed(164)
N=1000000
die= np.arange(6)+1
a=np.random.choice(die, N, replace=True)
np.count_nonzero(a==3)/N
▶ R
set.seed(1364)
N=1000000
a=sample(1:6, replace=T, N)
length(which(a==3))/N
▶ Matlab
rng(6356)
N=1000000;
a = randsample(6, N, true);
sum(a==3)/N
Probability
In general, the assignment of a probability to an event is rather
subtle
Sometimes there are some natural choices, for example based on
the existence of symmetries in the random experiment at hand
Sometimes the choice will be subjective (probability assignments
may differ from individual to individual)
Let us now consider a mathematical (axiomatic) definition of
probability:
probability as a function that satisfies some properties (a version of
these properties are indeed verified by the relative frequency)
Discrete Probability Space
Let the sample space Ω be non-empty and countable for the rest of
this chapter
Third element necessary to complete the description of a
probabilistic model is a function
P : Set of Events −→ [0, 1]
called probability that verifies the following two axioms
A1) P(Ω) = 1
The certain event has probability 1
A2) σ-additivity (countable additivity)
For every sequence (Ai)i∈N of disjoint/mutually exclusive
events, Ai ∩ Aj = ∅, i ̸= j
P
� ∞
�
i=1
Ai
�
=
∞
�
i=1
P(Ai)
(Ω, P) is called a discrete probability space, with Ω its sample
space, and subsets of Ω the events
Probability. Properties
For a discrete probability space (Ω, P), the following statements
hold true (all follow from A1 and A2)
S1) The impossible event has probability zero
P(∅) = 0
Proof. If Ai = ∅, i ∈ N, then �∞
i=1 Ai = ∅
P(∅) = P
� ∞
�
i=1
Ai
�
=
∞
�
i=1
P(Ai) =
∞
�
i=1
P(∅)
which holds iff P(∅) = 0.
Probability. Properties
S2) Finite additivity
Let Ai, i = 1, . . . , n, be a finite family of disjoint events,
Ai ∩ Aj = ∅, i ̸= j
P
� n�
i=1
Ai
�
=
n
�
i=1
P(Ai)
This follows from countable additivity, A2, setting Ak = ∅ for
all k ≥ n. Thus it is a weaker notion than countable additivity.
When Ω is finite, we can equally define the probability space
using axioms A1 and A2 or A1 and S2 (finite additivity)
Probability. Properties
Other consequences of A1 and A2 (draw the corresponding Venn
diagram if in doubt)
S3) P( ¯A) = 1 − P(A)
1 = P(Ω) = P(A ∪ ¯A) = P(A) + P( ¯A)
since A ∩ ¯A = ∅
S4) For any A, B ⊆ Ω
P(B \ A) = P(B) − P(A ∩ B)
Since B = (B \ A) ∪ (A ∩ B), with B \ A and (A ∩ B) disjoint,
then
P(B) = P(B \ A) + P(A ∩ B)
hence the result
Probability. Properties
S5) For any or any A, B ⊆ Ω
P(A ∪ B) = P(A) + P(B) − P(A ∩ B)
Since A ∪ B = A ∪ (B \ A), and A and B \ A are disjoint,
P(A ∪ B) = P(A) + P(B \ A) = P(A) + P(B) − P(A ∩ B)
using S4
Inclusion-Exclusion Formulae (Poincar´e’s Identities)
(Ω, P) a discrete probability space.
For any n ≥ 1 and for any choice of sets (events) A1, . . . An ⊆ Ω
P(A1 ∪ . . . ∪ An)
=
n
�
k=1
(−)k−1
�
1≤i1<...<ik≤n
P(Ai1 ∩ . . . ∩ Aik)
P(A1 ∩ . . . ∩ An)
=
n
�
k=1
(−)k−1
�
1≤i1<...<ik≤n
P(Ai1 ∪ . . . ∪ Aik)
They can be proven by induction
They are actually valid on any probability space (finite, countable
or uncountable)
Example
For example for n = 4, calling the events A1 = A, A2 = B,
A3 = C, A4 = D
P(A ∪ B ∪ C ∪ D)
=
P(A) + P(B) + P(C) + P(D)
−P(A ∩ B) − P(A ∩ C) − P(A ∩ D)
−P(B ∩ C) − P(B ∩ D) − P(C ∩ D)
+P(A ∩ B ∩ C) + P(A ∩ B ∩ D)
+P(A ∩ C ∩ D) + P(B ∩ C ∩ D)
−P(A ∩ B ∩ C ∩ D)
Alternating sum of the probabilities of each event (4 terms), each
possible pair of events (6), each possible triple of events (4), each
possible quadruple (1)
When to use the Poincar´e formula
Very often if one has to compute the probability that at least one
event occurs or the probability that no event occurs, the Poincar´e
formula is most useful
Indeed, let Ai be i = 1, . . . , n n events
P(A1 ∪ . . . ∪ An) is the probability that at least one of the n
events occurs
P( ¯A1 ∩ . . . ∩ ¯An) is the probability that none occurs
By the Poincar´e formulae, these probabilities can be written in
terms of probability involving fewer events, which are often easier
to compute
Probability of events and outcomes
(Ω, P) discrete probability space, A = {ωi1, . . . , ωik} a compound
event, then its probability is
P(A) =
k
�
j=1
P(ωij)
with the restriction 1 = P(Ω) = �
ω∈Ω P(ω)
Example
Ω = {ω1, ω2, ω3, ω4}
P(ωi) = p, ∀i
A1 = {ω1, ω2}, A2 = {ω1, ω3}, A3 = {ω1, ω4}
Find P(Ai), P(Ai ∩ Aj), P(A1 ∩ A2 ∩ A3), P(Ai ∪ Aj) etc.
Probability of events and outcomes
(Ω, P) discrete probability space, A = {ωi1, . . . , ωik} a compound
event, then its probability is
P(A) =
k
�
j=1
P(ωij)
with the restriction 1 = P(Ω) = �
ω∈Ω P(ω)
Example
Ω = {ω1, ω2, ω3, ω4}
P(ωi) = p, ∀i
A1 = {ω1, ω2}, A2 = {ω1, ω3}, A3 = {ω1, ω4}
Find P(Ai), P(Ai ∩ Aj), P(A1 ∩ A2 ∩ A3), P(Ai ∪ Aj) etc.
P(ωi) = 1/4, P(Ai) = 2/4, P(Ai ∩ Aj) = 1/4 = P(A1 ∩ A2 ∩ A3)
P(Ai ∪ Aj) = 3/4, P(A1 ∪ A2 ∪ A3) = P(Ω) = 1
Uniform Probability Space
A discrete probability space (Ω, P) which is finite (that is, Ω is
finite) and such that the outcomes ωi ∈ Ω are equiprobable is
called a uniform probability space.
In this case, then
P(ωi) = 1/|Ω|
The previous page example is such a space
More generally for any event E (a subset of Ω)
P(E) = |E|
|Ω|
The probability of an event is the ratio of the number of cases that
are favorable to it, to the number of possible cases, when there is
nothing to make us believe that one case should occur rather than
any other [Laplace]
Thus the problem of computing a probability of an event becomes
the problem of counting the number of its elements
Equiprobability
It is necessary that all points in the sample space be equiprobable
for computing probability via simple counting P(E) = |E|/|Ω|
Suppose we want to compute the probability of getting a three by
summing the numbers that turn up on tossing two dice
E = ”the sum of two throws is 3”
Since the sum is symmetric, we can think of using the following
sample space, where [i, j] is the unordered couple, with i the
number for one of the dice and j for the other
˜Ω =
[1, 1]
[1, 2]
[1, 3]
[1, 4]
[1, 5]
[1, 6]
[2, 2]
[2, 3]
[2, 4]
[2, 5]
[2, 6]
[3, 3]
[3, 4]
[3, 5]
[3, 6]
[4, 4]
[4, 5]
[4, 6]
[5, 5]
[5, 6]
[6, 6]
Equiprobability
While there may be nothing wrong in using ˜Ω, it should be noticed
however its simple events are not equally probable For example,
p([1, 1]) = 1/36
p([1, 2]) = p((1, 2) ∪ (2, 1)) = p((1, 2)) + p((2, 1)) = 2/36,
[i, j] unordered couple, (i, j) ordered couple
[i, j] one die shows the i-th face the other the j-th face
(i, j) the first die shows i the second j, if we throw them one after
the other or, if you toss them at the same time, just color the dice
differently: the red die shows i, the blue die j
[i, j] = {(i, j), (j, i)}
Thus we may not compute the probability by simple counting the
number of the points in the space that are favorable to the event
(one point [1, 2]) and divide by the size of the space 21 We would
get 1/21, instead of the correct probability which is 2/36
Equiprobability
Instead we can use the following equiprobable sample space
Ω =
(1, 1)
(1, 2)
(1, 3)
(1, 4)
(1, 5)
(1, 6)
(2, 1)
(2, 2)
(2, 3)
(2, 4)
(2, 5)
(2, 6)
(3, 1)
(3, 2)
(3, 3)
(3, 4)
(3, 5)
(3, 6)
(4, 1)
(4, 2)
(4, 3)
(4, 4)
(4, 5)
(4, 6)
(5, 1)
(5, 2)
(5, 3)
(5, 4)
(5, 5)
(5, 6)
(6, 1)
(6, 2)
(6, 3)
(6, 4)
(6, 5)
(6, 6)
whose |Ω| = 36 outcomes are all equiprobable P((i, j)) = 1/36,
thus |E = {i + j = 3}| = 2, P(E) = |E|/|Ω| = 2/36
More generally (exercise) the probability of rolling a sum of k, with
two dice is
P({i + j = k}) = 6 − |7 − k|
36
,
k = 2, . . . , 12
Combinatorics. Multiplication Rule
First rule of counting. Multiplication Rule
If an object is formed by making a succession of choices such that
there are n1 possibilities for the first choice, n2 possibilities for the
second (after the first choice is made) etc. then the total number
of objects that can be made by making a set of choices is
|E| = n1 · n2 · · ·
For the rule to apply, the number of available possibilities of each
choice must be the same irrespective of which choice was made
previously (ni for the i-th choice, which may be different from nj).
However the set of available possibilities may differ and depend on
the choice made at the previous stages
Book Problem 31
Beethoven wrote 9 symphonies, 5 piano concertos, and 32 piano
sonatas.
a) How many ways are there to play first a Beethoven symphony
and then a Beethoven piano concerto?
Book Problem 31
Beethoven wrote 9 symphonies, 5 piano concertos, and 32 piano
sonatas.
a) How many ways are there to play first a Beethoven symphony
and then a Beethoven piano concerto?
9 · 5 = 45
You may find it helpful to draw a tree
b) The manager of a radio station decides that on each
successive evening (7 days per week), a Beethoven symphony
will be played followed by a Beethoven piano concerto
followed by a Beethoven piano sonata. For how many years
could this policy be continued before exactly the same
program would have to be repeated?
Book Problem 31
Beethoven wrote 9 symphonies, 5 piano concertos, and 32 piano
sonatas.
a) How many ways are there to play first a Beethoven symphony
and then a Beethoven piano concerto?
9 · 5 = 45
You may find it helpful to draw a tree
b) The manager of a radio station decides that on each
successive evening (7 days per week), a Beethoven symphony
will be played followed by a Beethoven piano concerto
followed by a Beethoven piano sonata. For how many years
could this policy be continued before exactly the same
program would have to be repeated?
9 · 5 · 32 = 1440 days ≈ 4 years
Problem
How many ways can we roll three dice? In how many ways can
three dice appear when they are rolled? How many possible
numbers we get by rolling 3 dice (the order counts)?
We have a succession of three choices (one per die) and each die
represents a multiple choice of six possibilities ni = 6, thus there
are
6 · 6 · 6 = 216
possible rolls
This is obvious if you think of tossing the dice one after the other
rather than simultaneously (although it does not matter) The set
of choices at each throw of the dice in this case does not change
based on what happens at the previous stage (it is always one of
the numbers {1, 2, 3, 4, 5, 6} (and thus the number does not
change which is only thing that matters for the multiplicative rule)
Problem
In how many ways can two dice show different faces?
Problem
In how many ways can two dice show different faces?
6 · 5
In this case there are 6 available possibilities for the first choice.
Five different possibilities for the second die. Which five depends
on the first choice (e.g. if 3 shows up on the first roll, the set of
available possibilities is {1, 2, 4, 5, 6}, if 2 shows up, such set is
{1, 3, 4, 5, 6}), but there are always 5 possibilities for the second
die, so the first rule of counting as formulated above still applies
Problem
How many different calendars are possible for a year?
Problem
How many different calendars are possible for a year?
7 · 2 = 14
Each year starts on one of the seven days (Sunday, Monday, ...,
Saturday). Each year is either a leap year (i.e., it includes February
29) or not
Standard Ways of Counting and Basic Probabilistic Models
We are going to see now two standard ways of counting which can
be used in the majority of combinatorial problems
▶ Sampling Methods
▶ Allocation Methods
Basic Probabilistic Models. Sampling Model
Consider a population of n individuals (people, cards, numbered
balls), i.e. an aggregate of n distinguishable elements without
regard to their order (think of an urn containing nF numbered
balls)
Choose an individual from the population (a sample of size 1)
(that is, draw one ball from the urn)
How many ways can we do this?
Basic Probabilistic Models. Sampling Model
Consider a population of n individuals (people, cards, numbered
balls), i.e. an aggregate of n distinguishable elements without
regard to their order (think of an urn containing nF numbered
balls)
Choose an individual from the population (a sample of size 1)
(that is, draw one ball from the urn)
How many ways can we do this?
The answer is n
Choose k individuals successively, one at a time
How many ways can we do this?
Basic Probabilistic Models. Sampling Model
Consider a population of n individuals (people, cards, numbered
balls), i.e. an aggregate of n distinguishable elements without
regard to their order (think of an urn containing nF numbered
balls)
Choose an individual from the population (a sample of size 1)
(that is, draw one ball from the urn)
How many ways can we do this?
The answer is n
Choose k individuals successively, one at a time
How many ways can we do this?
It depends
Basic Probabilistic Models. Sampling Model
It depends on whether or not a chosen individual is returned to the
population before another is chosen (that is, on whether or not we
put the ball we have drawn back into the urn)
It also depends on whether the sample of size k is ordered or not
(namely, ordering means that the same individuals chosen in a
different order are considered to be a different sample- drawing 1
before 2 is considered different from drawing 2 before 1)
The two kinds of sampling models are called
sampling with replacement
we put the drawn ball back into the urn or we replace the ball
we have drawn in the urn
sampling without replacement
we do not replace the drawn balls
In addition we need to consider the sample to be ordered or
un-ordered
Basic Probabilistic Models
Thus we will consider the following 2 · 2 = 4 cases
▶ Sampling with replacement and with ordering
▶ Sampling without replacement and with ordering
▶ Sampling without replacement and without ordering
▶ Sampling with replacement and without ordering
Sampling with replacement and with ordering
I. Sampling with replacement and with ordering
The first individual is from a population with n individuals
n1 = n
We return the individual to the population, thus the set (and
hence the number) of possibilities for the choice of a second
individual is the same as for the first
n2 = n
Thus repeating the argument, using the multiplication rule (first
rule), we find that there are
n1 · · · nk = n · · · n = nk
ways to draw an (ordered) sample of size k from a population of n
individuals
Sampling with replacement and with ordering
nk counts the number of different ordered k-tuple from a
population of size n
Ordering means, e.g., that (1, 2, 1, 3) is considered different from
(1, 2, 3, 1), (2, 1, 3, 1), etc.
Put it differently, the sample space Ω is the set of ordered k-tuples
Ω = {(a1, a2, . . . , ak), ai = 1, ..., n}
Let us write down explicitly the number of ordered samples of size
k = 2 that can be obtained from n = 3 objects, if an object may
appear more than once (”sampling with replacement”)
Sampling with replacement and with ordering
nk counts the number of different ordered k-tuple from a
population of size n
Ordering means, e.g., that (1, 2, 1, 3) is considered different from
(1, 2, 3, 1), (2, 1, 3, 1), etc.
Put it differently, the sample space Ω is the set of ordered k-tuples
Ω = {(a1, a2, . . . , ak), ai = 1, ..., n}
Let us write down explicitly the number of ordered samples of size
k = 2 that can be obtained from n = 3 objects, if an object may
appear more than once (”sampling with replacement”)
(1, 1),
(1, 2),
(1, 3)
(2, 1),
(2, 2),
(2, 3)
(3, 1),
(3, 2),
(3, 3)